OpenStudy (anonymous):

Given below is the odd-number function. Which of the following are equal to O(30)? All that apply. O(n) = 2n - 1 A. The 30th odd number B. O(29)+1 C. 30 D. 59

OpenStudy (lilg132):

D

OpenStudy (lilg132):

dont forget to give a medal :)

OpenStudy (anonymous):

All that apply there are multiple right answers.

OpenStudy (anonymous):

O(30) = 2*30 - 1 = 59. a) 30th odd number = 59. Yes. b) O(29) = (2*29 - 1)+1 = 58. No. c) 30. No. d) 59. Yes. A and D.